# The long-term stability of the Vast Polar Structure and its connection to a possible previous passage of the LMC

### v1

@@ 3.2.1 Evolution of the VPOS from the MW on-plane satellites
We begin by analyzing the evolution of the VPOS using only the on-plane MW satellites. In Figure 3, we present the VPOS axis ratios $b/a$ and $c/a$, thickness, measured using both the RMS and the MAD, the direction of the normal vector in Galactic coordinates and its MRL, as a function of time. These metrics are shown for each of the six time-dependent gravitational potentials as lines, with shaded regions representing their associated uncertainties.
@@
The evolution of the axis ratios supports the long-term stability and planarity of the VPOS. Both $b/a$ and $c/a$ remain approximately constant across all gravitational potentials. In particular, $c/a$ is remarkably stable and consistently indicative of a flattened configuration. Its present-day values ($\sim 0.2$) are in agreement with previous VPOS characterizations (e.g., Pawlowski 2021) and remain consistent throughout the last 5 Gyr. We note that for potentials V23, L3M10, and L3M11, the values of $b/a$ and $c/a$ tend to be slightly lower at early times compared to the other potentials. This is likely due to Sculptor coming from large distances under these models, which elongates the satellite distribution and thus lowers the axis ratios. However, we note that these differences are small. The VPOS thickness also remains stable over time regardless of the potential. At $t=0$, both the RMS and MAD estimates yield a thickness of $\sim 15$ kpc, in line with previous studies. The RMS-based thickness shows mild fluctuations, occasionally reaching up to $\sim 35$ kpc, still within the range of thickness that current observational studies report. These variations likely reflect the complex orbital evolution of individual galaxies of the VPOS, particularly those with large apocentres, which can temporarily increase their projected distance to the plane, and thus increase its thickness. By contrast, the MAD measured thickness, being more robust to outliers, exhibits a remarkably flat trend, further backing that the VPOS has been a thin and persistent structure over time. The orientation of the VPOS exhibits some temporal variations, likely driven by the orbital motions of its member galaxies. As observed with the thickness, on-plane galaxies that momentarily increase their distance to the plane can shift the its orientation, leading to a temporary tilt. To better illustrate the directional stability of the VPOS, in Figure 4 we present an alternative visualization of the evolution of its orientation. The figure shows the Galactic longitude and latitude of the direction of the VPOS normal vector for potential V23 at each integration time, represented in a Cartesian $l$ vs. $b$ plot. While this is not an actual zoomed-in Galactic projection, it provides an approximation that allows us to visualize directional changes over time more clearly. The normal vector remains well within the VPOS area throughout the entire integration time, indicating that, despite some fluctuations, the overall orientation of the structure is stable. We note that results are very similar for the rest of potentials. The direction of the normal vector of the VPOS at $t=0$ ($l=182.2^{\circ};b=8.3^{\circ}$) differs from that reported by Pawlowski et al. (2013) in approximately 17 degrees, which is still well below the aperture defining the VPOS area ($\theta=36.87^{\circ}$). Such differences in the orientation of the VPOS between studies are likely attributable to the different galaxy samples used for its derivation. Finally, the MRL indicates that the distribution of normal vectors across the MC iterations is highly concentrated, with values remaining above $\sim 0.975$ throughout the whole integration time, highlighting the strong directional coherence of the structure.
@@
Uncertainties in the metrics tend to grow with look-back time due to the propagation of observational errors through the orbit integration process. However, this growth is not monotonic; uncertainties typically stabilize beyond $t\sim-1.5$ to $-2$ Gyr. Kumar et al. (2025) investigated how observational uncertainties affect the inferred stability of satellite planes using orbit integrations under MW-only potentials in a simulated VPOS. The study shows that while as $c/a$ may increase with time due to the observational uncertainties, this artificial thickening eventually plateaus or stabilizes after a few Gyr. In our study, we do not find significant increase in $c/a$ over time, and the inferred thickness remains low and stable across all potentials. Even if some level of artificial thickening is present, our results show that the VPOS remains a thin and coherent structure throughout the 5 Gyr integration period, reinforcing the evidence of its long-term persistence and stability.
@@
From $t=-2$ Gyr onward, we observe some subtle but potentially meaningful trends in the axis ratios and the direction of the normal vector. The $b/a$ ratio declines from $\sim$0.8 at $t\sim-1.5$ Gyr to $\sim$0.6 at tsimilar-$t\sim-$1 Gyr, suggesting an elongation of the satellite distribution within the plane, indicative of a more prolate configuration. Concurrently, the $c/a$ ratio exhibits a dip around $t\sim-0.5$ Gyr, indicating a transient vertical compression. This behavior is mirrored in the MAD thickness metric, which also reaches a local minimum at that epoch. Moreover, the orientation of the VPOS also shows changes over that time period. The reduced uncertainties at these recent times suggest that the changes are not a purely statistical fluctuation, but they are physically driven. Indeed, the timing of these structural changes is particularly notable, as they coincide with the recent infall of the LMC into the MW’s inner halo. Given the LMC’s well-established gravitational influence on the MW satellite system (see Vasiliev 2023), we interpret these changes in the VPOS structural parameters as a result of the dynamical response of the MW satellites to the LMC’s proximity. In fact, we find that, over the past $\sim$1 Gyr, the distribution of the MW on-plane galaxies tends to elongate towards the position of the LMC. The orientation of the longest principal axis measured for the distribution MW on-plane satellites also tends to align with the position of the LMC over the last $\sim$0.5 Gyrs. The orientation of the VPOS is particularly sensitive to such spatial rearrangements of its member galaxies. Although the LMC’s orbital path is well aligned with the VPOS, differing by only $\sim 5-15$ degrees prior to the LMC’s entry into the MW’s virial radius, even this modest offset induces subtle but measurable shifts in the VPOS normal vector. As a result of it, while the LMC approaches the VPOS from one side of the plane, its gravitational influence on the MW satellites alters their distribution and thus the VPOS orientation, tilting it, increasing both $l$ and $b$. Once the LMC trespasses the plane and moves away from it, the MW on-plane galaxies tend to rearrange accordingly, leading to the sudden change in the orientation of the VPOS at $t\sim-0.5$ Gyrs (notice the decrease of $l$ from $t\sim 0.5$ Gyrs onwards in Fig. 3 or the U-turn in the VPOS direction in Fig. 4). To sum up, we find that the LMC induces mild changes in the structural properties of the VPOS, without significantly altering its overall configuration or compromising its stability. This is likely a consequence of the close alignment between the LMC’s orbital path and the VPOS, which allows the otherwise intense gravitational influence of the LMC on the MW satellites during its recent approach to have a relatively smooth and coherent effect on the on-plane satellite population.
@@
Our results confirm that the VPOS has remained a stable structure throughout the entire integration time, regardless of the gravitational potential adopted for the orbit reconstruction. This allows us to place a lower limit on its age: the VPOS has persisted for at least 5 Gyr. Since our analysis so far considers only the MW satellites of the VPOS, we conclude that the planar structure was already in place around the MW well before the LMC crossed the virial radius. This rules out the possibility that the VPOS was formed as a result of the LMC’s recent infall. Interestingly, while the LMC had a dramatic dynamical impact on the broader MW satellite population in recent times, its effect on the VPOS has been relatively mild, not affecting its overall stability. This contrast raises important questions about the potential involvement of the LMC in the formation and evolution of the VPOS, especially given the good alignment between the LMC’s orbit and the VPOS plane.
@@ 3.3 Stability of the VPOS in the literature
Several observational studies have tried to assess the persistence of the VPOS using orbit integration, and unanimously conclude that it is a transient phenomenon (Lipnicky & Chakrabarti 2017; Maji et al. 2017; Sawala et al. 2023). Such analysis are focused on the 11 ’classical’ MW satellites (namely: Carina, Draco, Fornax, Leo I, Leo II, LMC, SMC, Sagittarius, Sculptor, Sextans, and Ursa Minor) and perform orbit integration under static MW-only potentials. Lipnicky & Chakrabarti (2017) performed backward integration using PMs from HST, finding that the VPOS goes from thickness 21.3 kpc at $t=0$ to disperse within 1 Gyr. They performed tests using several subsets of the classical satellites (whole sample, not considering Leo I and Leo II, and only with the best measured PMs), finding always the same fate for the VPOS. Maji et al. (2017) forward integrated the orbits of the classical dwarfs for a Gyr under 2 different MW potentials, finding that the VPOS members are going to abandon the planar structure. Their planar metrics $c/a$ and the thickness go from 0.18 and 19.6 kpc to 0.35 and 45 kpc at $t=-0.5$ Gyrs and finally 0.42 and 64 kpc at $t=-1$ Gyr. More recently Sawala et al. (2023), used PMs from Gaia EDR3 to forward and backward integrate the orbits of the classic satellites, finding that $c/a$ increases from $\sim 0.2$ to $\sim 0.3$ when advancing 0.5 Gyrs towards the future or 0.3 towards the past, concluding that the VPOS is transient.
@@
Our results show important differences with these studies. Although we obtain similar values of $c/a$ and thickness at present time, we do not find evidence of the VPOS vanishing, but the opposite. One possible explanation for these discrepancies is that previous studies are based on MW-only gravitational potentials for the orbit integration. However, the significant influence of the LMC on the orbits of MW satellites has been repeatedly demonstrated in recent years (Vasiliev 2023) and it is therefore essential to account for its effects when studying the stability of the VPOS. Gravitational potentials that include both the MW and the LMC provide a more realistic framework for capturing the complex orbital evolution of the galaxies in the vicinity of the MW. We note that the different modeling of the MW potential between previous studies and this work can also have an impact on the orbit reconstruction, nevertheless the inclusion of the effect of the LMC is likely to overshadow such differences. However, the most likely source of disagreement lies in the galaxy samples used to characterize the VPOS. The commonly studied set of 11 classical MW dwarf satellites presents some possible issues. While our analysis shows 7 of them to be clearly on-plane, Sagittarius and Sextans lie off-plane, and Leo I and Leo II are uncertain cases. Leo I and Leo II, despite being classified as ’uncertain’, can potentially modify the outcome of the analysis; not only are they located at large distances from the MW ($>200$ kpc, McConnachie 2012), but also their ambiguous membership to the VPOS is a result of a considerable scatter in their phase-space coordinates. The conjunction of these two factors, plus the inclusion of two off-plane galaxies in the analysis, can severely bias any stability or persistence assessment.
@@
In order to understand the origin of the discrepancies with previous studies, we repeated our analysis of the VPOS stability using as galaxy sample the classical satellites of the MW (see Appendix A). In Figure 6 we present the temporal evolution of key structural metrics of the plane defined by this subset of 11 galaxies. The axis ratios and thickness at $t=0$ support the presence of a planar configuration; however, this structure rapidly widens. The thickness increases quickly, particularly in the RMS-based measurement. The MAD-based thickness remains comparatively lower, reflecting its robustness to outliers and its better tracking of the stable, on-plane galaxies of the sample, though it also exhibits fluctuations. The orientation of the fitted plane also evolves substantially, deviating by more than 60 degrees from its initial direction and temporarily departing from the VPOS region before returning. In parallel, the MRL decreases with time, indicating that the normal vectors become increasingly dispersed across MC iterations. We conclude that the fast thickening and erratic orientation of the plane is driven by the off-plane and uncertain VPOS members, which introduce a significant scatter in the galaxy distribution and thus have a significant impact on the PCA analysis. The presence of on-plane galaxies in the sample of the classic satellites helps mitigate this effect to some extent, but it is not enough to keep the stability of the structure. Therefore, we conclude that the differences with previous studies regarding the stability and persistance of the VPOS are most likely due to the sample of galaxies used. This highlights the importance of carefully selecting the VPOS members, as including potential outliers has a dramatic impact on the study of its stability.
@@ 3.4 Origin and evolution of the VPOS
The analysis presented in this work shows that the VPOS is not a chance alignment in the present-day distribution of some dwarf galaxies around the MW, but rather a stable configuration that has persisted for at least the past 5 Gyr. This enduring stability raises questions about the underlying mechanisms that originated the VPOS. Understanding how this planar arrangement of galaxies formed and, more importantly, persisted over time could provide valuable insights into the history of the MW system and also help us understand how planes of satellites have formed elsewhere. The origin of the VPOS is a highly debated topic with no definitive solution yet (Pawlowski 2018). However, over the last years, several studies have helped to constrain the range of possible formation scenarios.
@@
Another scenario that has been extensively explored is the tidal dwarf galaxy (TDG) hypothesis, in which the VPOS would be formed from these second-generation galaxies created during a past major merger or close encounter between the MW and another galaxy. In this framework, TDGs arise from the tidal debris of such an interaction, forming along a common tidal tail and thus naturally sharing a common orbital plane and direction of motion (Pawlowski et al., 2011). However, a significant challenge for this scenario lies in the expected absence of dark matter in TDGs. As these systems are not embedded within dark matter halos, they would be expected to exhibit low mass-to-light ratios, an expectation that is inconsistent with observational evidence for MW satellites, which typically show mass-to-light ratios $\geq 10$ M/L (e.g., McConnachie 2012). However, these problems can be bypassed by adopting a series of alternative approaches like recurring cosmological frameworks different from $\Lambda$CDM, among others (see Pawlowski 2018, 2021). One advantage of this scenario is that it can be directly tested by studying the observational properties of the on-plane and off-plane galaxies, which are expected to be different given their separated origin. Taibi et al. (2024) did so by examining the properties of the galaxies of the MW-system, not finding a significant preference for on-plane and off-plane dwarfs to follow different scaling relations, more specifically by analysing their values on the luminosity-metallicity relation. Similar results have been reported for the plane around M31 (Collins et al. 2015). These results reduce the likelihood of the planar structures of the MW and M31 to be explained by TDGs.
Group infall has been proposed as another potential solution to explain the formation of planes of satellites (e.g. Li & Helmi 2008; D’Onghia & Lake 2008). The accretion of a group of dwarfs would result in the dispersion of its galaxies along a common plane around the host. Once the galaxies are accreted they would still share similar direction, energy and angular momentum. Recently, Taibi et al. (2024) studied the orbital histories of the MW on-plane systems finding that the co-orbiting ones are all currently moving towards the pericentres of their orbits, and had performed their apocentes $\sim 1$ Gyr ago. This striking similarity in the orbital phases of the on-plane MW satellites suggests that they may have originated from a group accretion event, in which multiple galaxies entered the MW halo together relatively recently.
The group infall hypothesis is particularly compelling in light of the ongoing accretion of the LMC system onto the MW. Having recently completed a pericentric passage, the LMC has significantly influenced the MW and its satellites throughout its trajectory across the Galactic halo (see Vasiliev 2023 for a review), and has also been proposed as a possible culprit of the formation of the VPOS. One proposed scenario is that the LMC could induce the clustering of the orbital poles of MW satellites (Garavito-Camargo et al. 2021); however, later studies argue that this effect alone cannot account for the formation of the VPOS (Correa Magnus & Vasiliev 2022; Pawlowski et al. 2022). Our analysis agrees that the recent passage of the LMC is not responsible for the formation of the VPOS, however we find that it reoriented the orbital pole of Grus II, bringing it into alignment with the structure. An alternative scenario posits that if the LMC is on its second pericentric passage, the VPOS could have naturally emerged from the stripping of some LMC satellites during the first pericentre (Vasiliev 2024). Although our study focuses on the stability of the VPOS and is agnostic regarding its origin, our findings consistently point to a possible link between the formation of the VPOS and the early orbital history of the LMC. We now explore the evidence and implications of such connection.

### v2

@@ 3.2.1 Evolution of the VPOS from the MW on-plane satellites
We begin by analyzing the evolution of the VPOS using only the on-plane MW satellites. In Figure 3, we present the VPOS axis ratios $b/a$ and $c/a$, thickness, measured using both the RMS and the MAD, the direction of the normal vector in Galactic coordinates and its MRL, as a function of time. These metrics are shown for each of the six potentials as lines, with shaded regions representing their associated uncertainties.
@@
The evolution of the axis ratios supports the long-term stability and planarity of the VPOS. Both $b/a$ and $c/a$ remain approximately constant across all gravitational potentials. In particular, $c/a$ is remarkably stable and consistently indicative of a flattened configuration. Its present-day values ($\sim 0.2$) are in agreement with previous VPOS characterizations (e.g., Pawlowski 2021b) and remain consistent throughout the last 5 Gyr. We note that for potentials V23, L3M10, and L3M11, the values of $b/a$ and $c/a$ tend to be slightly lower at early times compared to the other potentials. This is likely due to Sculptor coming from large distances under these models, which elongates the satellite distribution and thus lowers the axis ratios. However, we note that these differences are small. The VPOS thickness also remains stable over time regardless of the potential. At $t=0$, both the RMS and MAD estimates yield a thickness of $\sim 15$ kpc, in line with previous studies. The RMS-based thickness shows mild fluctuations, occasionally reaching up to $\sim 35$ kpc, still within the range of thickness that current observational studies report. These variations likely reflect the complex orbital evolution of individual galaxies of the VPOS, particularly those with large apocentres, which can temporarily increase their projected distance to the plane, and thus increase its thickness. By contrast, the MAD measured thickness, being more robust to outliers, exhibits a remarkably flat trend, further backing that the VPOS has been a thin and persistent structure over time. The orientation of the VPOS exhibits some temporal variations, likely driven by the orbital motions of its member galaxies. As observed with the thickness, on-plane galaxies that momentarily increase their distance to the plane can shift the orientation, leading to a temporary tilt. To better illustrate the directional stability of the VPOS, in Figure 4 we present an alternative visualization of the evolution of its orientation. The figure shows the Galactic longitude and latitude of the direction of the VPOS normal vector for potential V23 at each integration time, represented in a Cartesian $l$ vs. $b$ plot. While this is not an actual zoomed-in Galactic projection, it provides an approximation that allows us to better visualize directional changes over time. The normal vector remains well within the VPOS area throughout the entire integration time, indicating that, despite some fluctuations, the overall orientation of the structure is stable. We note that results are very similar for the rest of potentials. The direction of the normal vector of the VPOS at $t=0$ ($l=182.2^{\circ};b=8.3^{\circ}$) differs from that reported by Pawlowski et al. (2013) in approximately 17 degrees, which is still well below the aperture defining the VPOS area ($\theta=36.87^{\circ}$). Such differences in the orientation of the VPOS between studies are likely attributable to the different galaxy samples used for its derivation. Finally, the MRL indicates that the distribution of normal vectors across the MC iterations is highly concentrated, with values remaining above $\sim 0.975$ throughout the whole integration time, highlighting the strong directional coherence of the structure.
@@
Uncertainties in the metrics tend to grow with look-back time due to the propagation of observational errors through the orbit integration process. However, this growth is not monotonic; uncertainties typically stabilize beyond $t\sim-1.5$ to 2-2 Gyr. Kumar et al. (2025) investigated how observational uncertainties affect the inferred stability of satellite planes using orbit integrations under MW-only potentials in a simulated VPOS. The study shows that while as $c/a$ may increase with time due to the uncertainties, it eventually plateaus or stabilizes after a few Gyr. As also pointed out by Pawlowski (2021a), error sampling through a MC scheme, as done in this work, can be considered as applying uncertainties twice (first when measured, second when sampled), what can induce an artificial widening of the actual underlying plane when measuring it and thus lead to an underestimation of its stability. Given the mean uncertainty in the PMs ($\sim 0.03$ mas yr-1) and distance ($\sim 4\%$) for the MW on-plane galaxies, and based on the results of Kumar et al. (2025), we estimate that at t=3t=-3 Gyrs, the measured $c/a$ ratio could be $\sim 60-95\%$ larger than that of the underlying true plane. In our study, we do not find significant increase in $c/a$ over time, and the inferred thickness remains low and stable across all potentials. Even if some level of artificial thickening is present, our results show that the VPOS remains a thin and coherent structure throughout the 5 Gyr integration period, reinforcing the evidence of its long-term persistence and stability.
@@
From t=2t=-2 Gyr onward, we observe some subtle but potentially meaningful trends in the axis ratios and the direction of the normal vector. The $b/a$ ratio declines from $\sim$0.8 at $t\sim-1.5$ Gyr to $\sim$0.6 at $t\sim$-1 Gyr, suggesting an elongation of the satellite distribution within the plane, indicative of a more prolate configuration. Concurrently, the $c/a$ ratio exhibits a dip around $t\sim-0.5$ Gyr, indicating a transient vertical compression. This behavior is mirrored in the MAD thickness metric, which also reaches a local minimum at that epoch. Moreover, the orientation of the VPOS also shows changes over that time period. The reduced uncertainties at these recent times suggest that the changes are not a purely statistical fluctuation, but they are physically driven. Indeed, the timing of these changes is particularly notable, as they coincide with the recent infall of the LMC into the MW’s inner halo. Given the LMC’s well-established gravitational influence on the MW satellite system (see Vasiliev 2023), we interpret them as a result of the dynamical response of the MW satellites to the LMC’s proximity. In fact, we find that, over the past $\sim$1 Gyr, the distribution of the MW on-plane galaxies tends to elongate towards the position of the LMC. The orientation of the longest principal axis measured for the distribution MW on-plane satellites also tends to align with the position of the LMC over the last $\sim$0.5 Gyrs. However, we note that the dip in $c/a$ may arise from orbital phase effects: as most on-plane satellites are currently moving toward pericenter (Taibi et al., 2024), their past radial distribution would have been more extended, increasing the major axis while leaving the minor axis roughly unchanged. The orientation of the VPOS is particularly sensitive to spatial rearrangements of its member galaxies. Although the LMC’s orbital path is well aligned with the VPOS, differing by only $\sim 5-15$ degrees prior to the LMC’s entry into the MW’s virial radius, even this modest offset induces subtle but measurable shifts in the VPOS normal vector. As a result of it, while the LMC approaches the VPOS from one side of the plane, its gravitational influence on the MW satellites alters their distribution and thus the VPOS orientation, tilting it, increasing both $l$ and $b$. Once the LMC trespasses the plane and moves away from it, the MW on-plane galaxies tend to rearrange accordingly, leading to the sudden change in the orientation of the VPOS at $t\sim-0.5$ Gyrs (notice the decrease of $l$ from $t\sim 0.5$ Gyrs onwards in Fig. 3 or the U-turn in the VPOS direction in Fig. 4). To sum up, the LMC induces mild changes in the structural properties of the VPOS, without significantly altering its overall configuration or compromising its stability. This is likely a consequence of the close alignment between the LMC’s orbital path and the VPOS, which allows the otherwise intense gravitational influence of the LMC on the MW satellites during its recent approach to have a relatively smooth and coherent effect on the on-plane satellite population. This raises important questions about the potential involvement of the LMC in the formation and evolution of the VPOS.
@@
Our results confirm that the VPOS has remained a stable structure throughout the entire integration time, regardless of the gravitational potential adopted for the orbit reconstruction. This allows us to place a lower limit on its age: the VPOS has persisted for at least 5 Gyr. Since our analysis so far considers only the MW satellites of the VPOS, we conclude that the planar structure was already in place around the MW well before the LMC crossed the virial radius. This rules out the possibility that the VPOS was formed as a result of the LMC’s recent infall.
@@ 3.3 Stability of the VPOS in the literature
Several observational studies have tried to assess the persistence of the VPOS using orbit integration, and unanimously conclude that it is a transient phenomenon (Lipnicky & Chakrabarti 2017; Maji et al. 2017; Sawala et al. 2023). Such analysis are focused on the 11 ’classical’ MW satellites (namely: Carina, Draco, Fornax, Leo I, Leo II, LMC, SMC, Sagittarius, Sculptor, Sextans, and Ursa Minor) and perform orbit integration under static MW-only potentials. Lipnicky & Chakrabarti (2017) performed backward integration using PMs from HST, finding that the VPOS goes from thickness 21.3 kpc at $t=0$ to disperse within 1 Gyr. They performed tests using several subsets of the classical satellites (whole sample, not considering Leo I and Leo II, and only with the best measured PMs), finding always the same fate for the VPOS. Maji et al. (2017) forward integrated the orbits of the classical dwarfs for a Gyr under 2 different MW potentials, finding that the VPOS members are going to abandon the planar structure. Their planar metrics $c/a$ and the thickness go from 0.18 and 19.6 kpc to 0.35 and 45 kpc at t=0.5t=-0.5 Gyrs and finally 0.42 and 64 kpc at t=1t=-1 Gyr. More recently Sawala et al. (2023), used PMs from Gaia EDR3 to forward and backward integrate the orbits of the classic satellites, finding that $c/a$ increases from $\sim 0.2$ to $\sim 0.3$ when advancing 0.5 Gyrs towards the future or 0.3 towards the past, concluding that the VPOS is transient.
@@
Our results show important differences with these studies. Although we obtain similar values of $c/a$ and thickness at present time, we do not find evidence of the VPOS vanishing, but the opposite. One possible explanation for these discrepancies is that previous studies are based on MW-only gravitational potentials for the orbit integration. However, the significant influence of the LMC on the orbits of MW satellites has been repeatedly demonstrated in recent years (Vasiliev 2023) and it is therefore essential to account for its effects when studying the stability of the VPOS. Gravitational potentials that include both the MW and the LMC provide a more realistic framework for capturing the complex orbital evolution of the galaxies in the vicinity of the MW. We note that the different modeling of the MW potential between previous studies and this work can also have an impact on the orbit reconstruction, nevertheless the inclusion of the effect of the LMC is likely to overshadow such differences. However, the most likely source of disagreement lies in the galaxy samples used to characterize the VPOS. The commonly studied set of 11 classical MW dwarf satellites presents some possible issues. While our analysis shows 7 of them to be clearly on-plane, Sagittarius and Sextans lie off-plane, and Leo I and Leo II are uncertain cases. Leo I and Leo II, despite being classified as ’uncertain’, can potentially modify the outcome of the analysis; not only are they located at large distances from the MW (>200>200 kpc, McConnachie 2012), but also their ambiguous membership to the VPOS is a result of a considerable scatter in their phase-space coordinates, having both galaxies large uncertainties in their PMs and distances. The conjunction of these two factors, plus the inclusion of two off-plane galaxies in the analysis, can severely bias any stability or persistence assessment.
@@
In order to understand the origin of the discrepancies with previous studies, we repeated our analysis of the VPOS stability using as galaxy sample the classical satellites of the MW (see Appendix B). In Figure 6 we present the temporal evolution of key structural metrics of the plane defined by this subset of 11 galaxies. The axis ratios and thickness at $t=0$ support the presence of a planar configuration; however, this structure rapidly widens. The thickness increases quickly, particularly in the RMS-based measurement. The MAD-based thickness remains comparatively lower, reflecting its robustness to outliers and its better tracking of the stable, on-plane galaxies of the sample, though it also exhibits fluctuations. The orientation of the fitted plane also evolves substantially, deviating by more than 60 degrees from its initial direction and temporarily departing from the VPOS region before returning. In parallel, the MRL decreases with time, indicating that the normal vectors become increasingly dispersed across MC iterations. We conclude that the fast thickening and erratic orientation of the plane is driven by the off-plane and uncertain VPOS members, which introduce a significant scatter in the galaxy distribution and thus have a significant impact on the PCA analysis. We note that Leo I and Leo II are the main contributors to the rapid widening of the plane, given that their large uncertainties in PMs and distances lead to poorly constrained orbits that strongly influence the plane fit. The presence of on-plane galaxies in the sample of the classic satellites helps mitigate this effect to some extent, but it is not enough to keep the stability of the structure. Therefore, we conclude that the differences with previous studies regarding the stability and persistance of the VPOS are most likely due to the sample of galaxies used. This highlights the importance of carefully selecting the VPOS members, as including potential outliers has a dramatic impact on the study of its stability.
@@ 3.4 Origin and evolution of the VPOS
The analysis presented in this work shows that the VPOS is not a chance alignment in the present-day distribution of some dwarf galaxies around the MW. Understanding how this planar arrangement of galaxies formed and, more importantly, persisted over time could provide valuable insights into the history of the MW system and also help us understand how planes of satellites have formed elsewhere. The origin of the VPOS is a highly debated topic with no definitive solution yet (Pawlowski 2018). However, over the last years, several studies have helped to constrain the range of possible formation scenarios.
@@
Group infall has been proposed as another potential solution to explain the formation of planes of satellites (e.g. Li & Helmi 2008; D’Onghia & Lake 2008; Jerjen et al. 2025). The accretion of a group of dwarfs would result in the dispersion of its galaxies along a common plane around the host. Once the galaxies are accreted they would still share similar direction, energy and angular momentum. Recently, it has also been explored the role of group infall in shaping the MW stellar halo (Callingham & Helmi 2025), and potential evidence of a group accretion event has been reported among the MW satellites (Leo II, Leo IV, and Crater 1 , Júlio et al. 2024). Also recently, Taibi et al. (2024) studied the orbital histories of the MW on-plane systems finding that the co-orbiting ones are all currently moving towards the pericentres of their orbits, and had performed their apocentes $\sim 1$ Gyr ago. This striking similarity in the orbital phases of the on-plane MW satellites suggests that they may have originated from a group accretion event, in which multiple galaxies entered the MW halo together relatively recently.
The group infall hypothesis is particularly compelling in light of the ongoing accretion of the LMC system onto the MW. Having recently completed a pericentric passage, the LMC has significantly influenced the MW and its satellites throughout its trajectory across the Galactic halo (see Vasiliev 2023 for a review), and has also been proposed as a possible culprit of the formation of the VPOS. One proposed scenario is that the LMC could induce the clustering of the orbital poles of MW satellites (Garavito-Camargo et al. 2021); however, later studies argue that this effect alone cannot account for the formation of the VPOS (Correa Magnus & Vasiliev 2022; Pawlowski et al. 2022). Our analysis agrees that the recent passage of the LMC is not responsible for the formation of the VPOS, however we find that it reoriented the orbital pole of Grus II, bringing it into alignment with the structure. An alternative scenario posits that if the LMC is on its second pericentric passage, the VPOS could have naturally emerged from the stripping of some LMC satellites during the first pericentre (Vasiliev 2024). Previous works had also proposed a possible association between some of the on-plane galaxies and the LMC. Lynden-Bell (1976) suggested that the galaxies known at the time to lie in the plane could have been tidally stripped from the “Great Magellanic Galaxy,” the alleged progenitor of the LMC. Some subsequent studies similarly argued that several of the classical MW satellites may have once been associated with the LMC (Nichols et al. 2011) and some remarked the close connection between the VPOS, the LMC, and the Magellanic Stream (Pawlowski et al. 2013, 2015). Although our study focuses on the stability of the VPOS and is agnostic regarding its origin, our findings consistently point to a possible link between the formation of the VPOS and the early orbital history of the LMC. We now explore the evidence and implications of such connection.
3.5 The LMC and the origin of the VPOS
First, we note that the orbital path of the LMC aligns well with the VPOS (see Fig. 2, and Section 3.2.1). If the LMC is indeed on its first infall into the MW halo (Besla et al. 2007), such chance alignment would be highly improbable (Metz et al. 2008). The odds of an external group arriving from large distances along a trajectory that precisely traverses a long-lasting planar structure around the MW are very low, assuming they are completely unrelated. This apparent coincidence becomes even more surprising if we consider that the direction of motion of the LMC system around the MW also aligns with the sense of rotation of the VPOS MW satellites. To quantify this, we estimated the probability that two randomly oriented vectors are separated by an angle less than $\theta$, taken here to be between $\sim 5^{\circ}$ and $15^{\circ}$, consistent with the offset between the LMC orbit and the VPOS prior to the LMC entering the MW’s virial radius. This corresponds to the solid angle of a spherical cap of angular radius $\theta$, normalized by the total surface area of a sphere, yielding a probability of approximately $\sim 0.2$$1.7\%$. Including the additional condition that the orbital sense must also match further reduces this probability by half. Altogether, this suggests that the LMC and its satellites are not merely recent additions to the planar structure around the MW.